# HAFixAgent: History-Aware Automated Program Repair Agent

### v2

@@ 7.1. Performance Regressions
We observe 32 cases where a history configuration underperforms the non-history one or misses non-history fixes (Figure 3). For instance, in MFMH, the non-history configuration attains 50 while fn_all, fn_pair, and fl_diff reach 47, 43, and 48, respectively, even though each contributes 12 to 13 unique fixes that the non-history does not solve. This pattern aligns with LLM evidence that more context can hurt: models are easily distracted by irrelevant additions (shi2023large), and performance often declines when relevant facts sit in the middle or deeper parts of long prompts (levy2024same). In our setting, extra historical snippets can dilute salient context and reduce Plausible@1 on some bugs that don’t need historical context. Future work might filter and summarize history for direct relevance or add a guardrail that falls back to a leaner prompt when early steps show no progress.
7.2. Heuristics Complementarity
A key finding from RQ1 (Figure 3) is that the three historical heuristics are complementary, not redundant. Each historical heuristic (e.g., fn_all, fn_pair, fl_diff) repairs a distinct set of bugs that the others miss. This has important implications for why history is effective. It suggests that different bug types are susceptible to different forms of historical context:
  • fl_diff (file-level diff), the strongest overall, likely excels at capturing fine-grained and localized edit patterns that are common in single-line and single-hunk bugs.
  • fn_pair (historical function before and after) provides a richer semantic context, showing the agent how a function’s logic was intended to operate, which may be crucial for more complex logic errors.
  • fn_all (all functions from blame commit) captures co-evolutionary patterns (e.g., when function A changed, function B also changed), which is a vital clue for multi-hunk bugs where a fix requires coordinated edits.
This complementarity also explains the practical trade-off identified in RQ2 (Figure 6): combining two or three heuristics shows a higher pass rate and yields the most cost-effective performance, as it provides the agent with a diverse set of contexts.
7.3. Future Work
Richer historical heuristics beyond the blamed commit.
HAFixAgent currently injects three compact heuristics drawn from the single commit that last touched the buggy lines. A broader temporal view may help: commits immediately before or after the blamed change, refactorings, and coevolution bursts can reveal the mechanism that produced the bug. Future work can build coevolution-aware retrieval over historical commit windows, summarize change intents such as API migrations or dependency upgrades, and learn when to expand beyond one commit. Such a broader historical context may improve robustness, particularly for bugs whose causes span multiple commits.
Enhanced context for blameless bugs.
Building on our discussion in Section 4.1.3, the simple nearest-line fallback for the 28.9% of blameless bugs can be significantly improved. Instead of just finding the nearest line, a future agent could analyze the history of the parent class or module to infer project-specific patterns for adding new functionality. It could also search the repository for semantically similar add-only commits (e.g., how has this project added null-checks in other files?) to provide the agent with a template for the kind of code it needs to write.
Assessing history’s impact on patch quality.
A valuable future direction is to move beyond test-pass-based evaluation to investigate how historical context impacts the quality and maintainability of the generated patches. While multiple agent configurations might produce a plausible fix, it is an open question whether the history-aware patches are qualitatively superior. For instance, a patch generated with historical context might be more idiomatic, such as re-using a specific helper function or logging style from the historical diff, and more semantically aligned with the project’s evolution. This could contrast with a non-history patch that, while functional, acts as a brute-force fix that simply passes the test suite. Future work could investigate this by designing automated tools to rate patch quality or by employing static analysis metrics to measure code complexity and semantic differences.
Extending to a multi-agent repair system.
A promising, though more complex, frontier is to evolve HAFixAgent from a single-agent system into a collaborative multi-agent architecture (rafi2024multi; DBLP:conf/icse/LinKC25; lee2024unified). Instead of one agent handling all tasks, we could design a system with specialized roles. For instance, a Coordinator agent could oversee the process, dispatching a specialized History agent to analyze repository patterns, blame commits, and issue reports to form a repair hypothesis. This hypothesis would then be passed to a Fixer agent focused on patch generation. Critically, a Reviewer agent could then critique the proposed patch, checking it not only against the test suite but also against the historical context for style, conventions, and potential side effects, which is similar to human code review. This separation of concerns could be particularly effective for the complex multi-file-multi-hunk (MFMH) bugs, where a single agent’s reasoning chain and context window can be overwhelmed.
Applying historical context to other SE tasks beyond APR
The same history-based heuristics may benefit other SE code tasks such as code generation, code completion, code review, refactoring, and test generation. Repository history heuristics such as coevolution patterns, change intent, and dependency updates can guide model choices and keep outputs aligned with project conventions. Future work should evaluate these heuristics on tasks beyond repair using task-specific metrics (functional correctness, style consistency, temporal consistency, and safety), compare against retrieval baselines, and study generalization across languages and repositories.
@@ 8. Threats to Validity
In this section, we discuss the threats to validity of our study about HAFixAgent.
8.1. Internal Validity
Data leakage.
@@ 8.1. Internal Validity
DeepSeek-V3.2-Exp is available as open weights, but the pretraining corpus for this model is not disclosed to the best of our knowledge. As a result, we cannot verify if the specific projects or bugs in Defects4J are included in the pretraining data. However, our evaluation of HAFixAgent focuses on the relative effectiveness improvement of different historical heuristics, rather than their absolute performance. For instance, if HAFixAgent-fl_diff fixes a bug that the HAFixAgent-non-history does not fix, this improvement is attributed to the benefit of historical context, regardless of whether the LLM has seen the bug before. Our evaluation design naturally isolates and mitigates the possible influence of potential data leakage. To fully resolve this threat, we would need to retrain DeepSeek-V3.2-Exp from scratch, which would be infeasible for an academic project.
Fault localization.
Our evaluation assumes perfect fault localization to isolate the evaluation of historical context on patch generation, a common setup in LLM-based APR studies (yang2024sweagent; fan2023automated; bouzenia2024repairagent; xia2022less; ye2024iter; nashid2025characterizing; ehsani2025bug). Note that even given perfect fault localization, as most of the current work did, the LLM-based APR still struggle with complicated bugs (e.g., 39.1% Plausible for non-history configuration in MFMH bugs in our study). In practice, FL is noisy, especially for human-reported issues and multi-hunk defects, which can lower repair rates and lead to potentially wrong repair directions. For iterative agentic APR, updating FL during repair might help, but it does not eliminate the challenge. So we leverage perfect fault localization to isolate the evaluation of the historical context in HAFixAgent.
8.2. External Validity
Dataset and language generalizability.
@@ 8.2. External Validity
Our study uses Defects4J v3.0.1 with 854 Java bugs. Results may not transfer to other ecosystems where language, project size, test adequacy, and repository practices differ. The history profile we observe in Defects4J (for example, many bugs mapping to a single history commit) can change in other benchmarks, repositories with frequent code moves, or workflows that rewrite history. To mitigate this threat, we expose a dataset-agnostic interface behind the history extractor and provide dataset adapters that can be subclassed when porting to languages such as Python or C. We release code and scripts so other communities can rerun the pipeline on local datasets.
LLM and agent generalizability.
@@
HAFixAgent is instantiated with one agent loop, a narrow bash tool set, specific prompts, and a single LLM configuration with fixed step, cost, and time guards. Different agents, tool scopes, models, context windows, or decoding policies may change absolute numbers and the relative strength of the three history heuristics. API level determinism does not guarantee identical runs in practice, so small variation across executions and providers is expected. To reduce this threat, we use DeepSeek-V3.2-Exp as a recently released strong model at the time of writing. We also assume perfect fault localization to isolate the effect of history, and generalizing to noisy localization remains future work.
@@ 9.1. Traditional Automated Program Repair
Traditional APR formulates patch generation as a search or constraint problem over program transformations guided by tests or specifications. Search-based systems evolve patches with genetic or heuristic search over mutation operators and validate candidates against the test suite, exemplified by GenProg and successors (GenProg; le2016history; wen2018context). Semantics or constraint-based repair derives patches by solving synthesis constraints obtained from symbolic execution or program analysis, as in SemFix (nguyen2013semfix) and Angelix (mechtaev2016angelix), and with targeted condition synthesis such as Nopol (xuan2016nopol). Template-based repair uses human-designed or mined fix schema that match buggy contexts and instantiate edits such as TBar (liu2019tbar) and AVATAR (liu2019avatar). A complementary line reuses in-project code fragments as ingredients, such as ssFix (DBLP:conf/kbse/XinR17) and CAPGEN (wen2018context). These approaches established core pipelines for localization, candidate generation, and test-based validation, but their coverage is constrained by operator or pattern design and by search scalability.
Learning-based APR before LLMs frames repair as code-to-code translation or edit prediction trained on a large amount of code corpus. Early neural systems include DeepFix (gupta2017deepfix) for syntax errors and sequence-to-sequence repair, such as SequenceR (chen2019sequencer), with later improvements from model ensembling and syntax guidance in CoCoNuT (lutellier2020coconut), CURE (jiang2021cure), and Recoder (zhu2021syntax). Template mining and ranking were also brought to production, for example, Facebook’s Getafix (bader2019getafix) for static-analyzer warnings and the end-to-end SapFix (marginean2019sapfix) pipeline that proposed developer-reviewed patches. Together, traditional APR demonstrates effective repair across bug classes while revealing recurring limitations in computation cost, low accuracy and overfitting to tests. These limitations have motivated LLM and agent-based APR that learn fixes from data and plan tool-driven repair loops, which we introduce next.
@@ 9.2. LLM and Agent based Automated Program Repair
Large language models (LLMs) have significantly advanced Automated Program Repair (APR), moving from direct prompting to sophisticated, multi-step agentic systems. Early evidence shows code LLMs already outperform prior learning-based APR and benefit from fine-tuning and careful evaluation on standard benchmarks (jiang2023impact; xia2023automated). This led to a variety of prompting and interaction strategies. For example, conversation-driven repair simulates a developer’s debugging process by interleaving patch generation with test feedback, achieving a strong balance of cost and accuracy (xia2024automated). Other approaches focus on improving the model’s reasoning, such as self-directed repair, which uses a chain-of-thought process to gather knowledge before fixing (yin2024thinkrepair). The plastic surgery line revisits ingredient reuse by aligning LLMs with project-specific code (xia2023plastic). Beyond pure prompting, hybrid and template-guided methods constrain or scaffold the patch space with analysis or repair templates to improve plausibility and correctness (li2024hybrid; huang2025template).
Agent-based APR systems represent a further step, shifting from single-shot patch generation to autonomous, tool-using workflows. Pioneering general-purpose software agents, like SWE-agent (yang2024sweagent) and OpenHands (wang2024openhands), established a powerful paradigm by equipping LLMs with tools to interact with a repository, such as editors, shell commands, and test runners. Building on this, repair-focused agents have integrated more specialized tools. For instance, some employ search and fault localization to narrow down the buggy code (bouzenia2024repairagent; zhang2024autocoderover), while others explore multi-agent collaboration to divide the debugging task (lee2024unified), with empirical studies beginning to map this rapidly evolving design space (meng2024empirical). A recent trend also focuses on improving agent performance on complex, repository-level tasks by incorporating memory and experience, allowing agents to learn from prior repair trajectories (mu2025experepair; chen2025swe). These lines generally rely on local code and runtime outputs, with limited study of commit history as a first-class context inside the loop, which is the focus of our work.
@@ 9.3. In-context Learning for Automated Program Repair
A core question for LLM-based APR is which context to provide and how to structure it. Prior work shows that adding bug-related facts such as error messages, stack traces, and failing tests improves repair, and that the choice of facts matters (parasaram2024fact; xia2024automated). Studies on local context indicate that models are sensitive to how much and which code surrounds the edit region (prenner2024out). Other work leverages natural language bug reports or transforms descriptions into edits, and explores execution traces to guide patching (fakhoury2024nl2fix; haque2025towards). Repository-level evaluations report that larger, realistic tasks remain challenging and require broader context curation (chen2024large).
@@
Beyond local signals, historical context is receiving attention. HAFix (shi2025hafix) demonstrates that blame-derived commit data is an effective, lightweight historical signal for fixing more bugs, motivating history-aware agent designs at Defects4J scale. Layered knowledge injection systematically adds project and repository knowledge to prompts and improves bug fixing on broader-context scenarios (ehsani2025bug). Built on top of HAFix, HAFixAgent advances this direction by bringing blame-driven history into the agentic workflow and quantifying its effect on single-hunk and multi-hunk bugs.
10. Conclusion
@@
This paper presented HAFixAgent, a history-aware agentic approach that injects blame-derived repository context into an observe-act-verify loop to address the more complex single- and multi-hunk bug categories. A preliminary study shows how blame commit history is not only widely available (71.1% blameable), but also is surprisingly concentrated with 70.7% of Defects4J bugs having exactly one unique blame commit across their buggy line.
When applying HaFixAgent on all 854 Defects4J bugs, history-aware settings were able to add 194 history-only fixes that the non-history configuration never achieves. When compared to SOTA baselines, HAFixAgent outperforms RepairAgent by an average of 212.3% and BIRCH feedback by 29.9%. Among the three historical contexts, fl_diff is the strongest overall and fn_pair leads on single-hunk bugs. Successful runs finish well before the 50-step cap, and the median token cost does not increase; in the hardest multi-file-multi-hunk category, costs are lower than the non-history configuration. Lastly, combining two or three heuristics captures most gains, while leveraging all heuristics adds little benefit for a significantly higher cost. These findings give a simple recipe for practice: (1) ground the agent loop in version control history, (2) prefer diff-based context, (3) add one complementary heuristic when needed, and (4) avoid collecting excessive context.
This study has natural limitations. We assume perfect fault localization (similar to related work), focus on one benchmark and language, and use a single strong agent model and loop. Future work will relax these assumptions by evaluating additional datasets and languages, testing imperfect localization, learning to select or combine history heuristics at runtime, and integrating richer project artifacts. We expect progressive use of historical context to remain a robust and lightweight direction for advancing agent-based program repair.

### v3

@@ 7. Discussion
Performance Regressions. We observe 32 cases where a history configuration underperforms the non-history one or misses non-history fixes (Figure 3). For instance, in MFMH, the non-history configuration attains 50 while fn_all, fn_pair, and fl_diff reach 47, 43, and 48, respectively, even though each contributes 12 to 13 unique fixes that the non-history does not solve. This pattern aligns with LLM evidence that more context can hurt: models are easily distracted by irrelevant additions (Shi et al., 2023), and performance often declines when relevant facts sit in the middle or deeper parts of long prompts (Levy et al., 2024). In our setting, extra historical snippets can dilute salient context and reduce Plausible@1 on some bugs that don’t need historical context.
Heuristics Complementarity. A key finding from RQ1 (Figure 3) is that the three historical heuristics are complementary, not redundant, with each repairing a distinct set of bugs. This suggests that different bug types are susceptible to different forms of historical context: fl_diff excels at fine-grained edit patterns (SL/SH), fn_pair provides semantic context showing function evolution (complex logic errors), and fn_all captures co-evolutionary patterns across functions (multi-hunk bugs). The strongest benefits appear for SFMH bugs, which involve multiple related changes within a single file. Because the blame commit typically captures the entire change context for such bugs (as shown by the single-commit concentration in RQ0), all three heuristics provide highly relevant information about the bug-introducing pattern. This complementarity also explains the practical trade-off identified in RQ2: combining two or three heuristics yields the most cost-effective performance.
Plausible vs. Correct Patches. Our evaluation uses plausible patches (passing the full test suite) as the success metric, consistent with prior work (Nashid et al., 2025; Ehsani et al., 2025; Silva and Monperrus, 2025). Manual verification of over 3.5k plausible patches (4 configurations $\times$ 2 datasets) is prohibitively expensive at this scale. Our evaluation focuses on relative improvement rather than absolute correctness: the same plausible metric and same LLM are used consistently across all comparisons, including both replicated baselines and the history vs. non-history ablation, ensuring fair comparison. More broadly, as LLM and agent-based APR systems generate plausible patches at unprecedented scale, designing automated patch correctness validation becomes a pressing open problem for the community.
@@ 8. Threats to Validity
Internal Validity. DeepSeek-V3.2-Exp is available as open weights, but the pretraining corpus is not disclosed to our knowledge. As a result, we cannot verify if the specific projects or bugs in Defects4J are included in the pretraining data. However, our evaluation focuses on the relative effectiveness improvement of different historical heuristics, rather than their absolute performance. For instance, if HAFixAgent-fl_diff fixes a bug that HAFixAgent-non-history does not fix, this improvement is attributed to the benefit of historical context, regardless of whether the LLM has seen the bug before. Furthermore, the substantially larger history benefit on BugsInPy (+38.6%) compared to Defects4J (+4.4%) supports this interpretation: if data leakage explained the gains, both benchmarks should show similar history benefits, since the same LLM and same ablation design are used. The larger BugsInPy effect suggests that history provides genuine signal that is more impactful when the LLM has less memorized knowledge about the benchmark.
Our main evaluation assumes perfect fault localization, a common setup in LLM-based APR studies (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025). To mitigate this threat, RQ2a evaluates robustness under noisy FL (line shifts of +1/+3/+5), showing that history-aware configurations maintain effectiveness even when FL is imprecise. However, our noise model (uniform positive shifts) does not capture all real-world FL errors, such as entirely wrong file predictions.
@@
External Validity Our study evaluates on Defects4J (854 Java bugs) and BugsInPy (501 Python bugs). While this covers two languages and 34 projects, results may not transfer to other ecosystems where project size, test adequacy, and repository practices differ. The history profiles already differ between our two benchmarks (e.g., 99.5% single-commit in Defects4J vs. 58.4% in BugsInPy), suggesting that history-aware techniques may need adaptation per ecosystem. We expose a dataset-agnostic interface behind the history extractor and release all code and scripts for replication. Evaluating on repository-level benchmarks such as SWE-bench (Jimenez et al., 2024) is future work.
@@
HAFixAgent is instantiated with one agent loop, a narrow bash tool set, specific prompts, and a single LLM with fixed step, cost, and time guards. Different agents, models, or decoding policies may change absolute numbers; API-level determinism does not guarantee identical runs. To reduce this threat, we use DeepSeek-V3.2-Exp and replicate all baselines on the same LLM.
@@ 9.1. Traditional Automated Program Repair
Traditional APR formulates patch generation as a search or constraint problem over program transformations guided by tests or specifications. Search-based systems evolve patches with genetic or heuristic search over mutation operators and validate candidates against the test suite, exemplified by GenProg and successors (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018). Semantics or constraint-based repair derives patches by solving synthesis constraints obtained from symbolic execution or program analysis, as in SemFix (Nguyen et al., 2013) and Angelix (Mechtaev et al., 2016), and with targeted condition synthesis such as Nopol (Xuan et al., 2017). Template-based repair uses human-designed or mined fix schema that match buggy contexts and instantiate edits such as TBar (Liu et al., 2019b) and AVATAR (Liu et al., 2019a). A complementary line reuses in-project code fragments as ingredients, such as ssFix (Xin and Reiss, 2017) and CAPGEN (Wen et al., 2018). These approaches established core pipelines for localization, candidate generation, and test-based validation, but their coverage is constrained by operator or pattern design and by search scalability.
Learning-based APR before LLMs frames repair as code-to-code translation or edit prediction trained on a large amount of code corpus. Early neural systems include DeepFix (Gupta et al., 2017) for syntax errors and sequence-to-sequence repair, such as SequenceR (Chen et al., 2021b), with later improvements from model ensembling and syntax guidance in CoCoNuT (Lutellier et al., 2020), CURE (Jiang et al., 2021), and Recoder (Zhu et al., 2021). Template mining and ranking were also brought to production, for example, Facebook’s Getafix (Bader et al., 2019) for static-analyzer warnings and the end-to-end SapFix (Marginean et al., 2019) pipeline that proposed developer-reviewed patches. Together, traditional APR demonstrates effective repair across bug classes while revealing recurring limitations in computation cost, low accuracy and overfitting to tests. These limitations have motivated LLM and agent-based APR that learn fixes from data and plan tool-driven repair loops, which we introduce next.
@@ 9.2. LLM and Agent based Automated Program Repair
Large language models (LLMs) have significantly advanced Automated Program Repair (APR), moving from direct prompting to sophisticated, multi-step agentic systems. Early evidence shows code LLMs already outperform prior learning-based APR and benefit from fine-tuning and careful evaluation on standard benchmarks (Jiang et al., 2023; Xia et al., 2023b). This led to a variety of prompting and interaction strategies. For example, conversation-driven repair simulates a developer’s debugging process by interleaving patch generation with test feedback, achieving a strong balance of cost and accuracy (Xia and Zhang, 2024). Other approaches focus on improving the model’s reasoning, such as self-directed repair, which uses a chain-of-thought process to gather knowledge before fixing (Yin et al., 2024). The plastic surgery line revisits ingredient reuse by aligning LLMs with project-specific code (Xia et al., 2023a). Beyond pure prompting, hybrid and template-guided methods constrain or scaffold the patch space with analysis or repair templates to improve plausibility and correctness (Li et al., 2025a; Huang et al., 2025).
Agent-based APR systems represent a further step, shifting from single-shot patch generation to autonomous, tool-using workflows. Pioneering general-purpose software agents, like SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025), established a powerful paradigm by equipping LLMs with tools to interact with a repository, such as editors, shell commands, and test runners. Building on this, repair-focused agents have integrated more specialized tools. For instance, some employ search and fault localization to narrow down the buggy code (Bouzenia et al., 2025; Zhang et al., 2024c), while others explore multi-agent collaboration to divide the debugging task (Lee et al., 2024), with empirical studies beginning to map this rapidly evolving design space (Meng et al., 2024). A recent trend also focuses on improving agent performance on complex, repository-level tasks by incorporating memory and experience, allowing agents to learn from prior repair trajectories (Mu et al., 2025; Chen et al., 2025). These lines generally rely on local code and runtime outputs, with limited study of commit history as a first-class context inside the loop, which is the focus of our work.
@@ 9.3. In-context Learning for Automated Program Repair
A core question for LLM-based APR is which context to provide and how to structure it. Prior work shows that adding bug-related facts such as error messages, stack traces, and failing tests improves repair, and that the choice of facts matters (Parasaram et al., 2025; Xia and Zhang, 2024). Studies on local context indicate that models are sensitive to how much and which code surrounds the edit region (Prenner and Robbes, 2024). Other work leverages natural language bug reports or transforms descriptions into edits, and explores execution traces to guide patching (Fakhoury et al., 2024; Haque et al., 2025). Repository-level evaluations report that larger, realistic tasks remain challenging and require broader context curation (Chen et al., 2024).
@@
Beyond local signals, historical context is receiving attention. HAFix (Shi et al., 2025) demonstrates that blame-derived commit data is an effective, lightweight historical signal for fixing more bugs, motivating history-aware agent designs at Defects4J scale. Layered knowledge injection systematically adds project and repository knowledge to prompts and improves bug fixing on broader-context scenarios (Ehsani et al., 2025). Built on top of HAFix, HAFixAgent advances this direction by bringing blame-driven history into the agentic workflow and quantifying its effect on single-hunk and multi-hunk bugs.
10. Conclusion and Future Work
@@
This paper presented HAFixAgent, a history-aware agentic approach that injects blame-derived repository context into an observe-act-verify loop to address the more complex single- and multi-hunk bug categories. A preliminary study on 854 Defects4J (Java) and 501 BugsInPy (Python) bugs shows that blame commit history is widely available (71.1% and 87.4% blameable, respectively), with Defects4J blame highly concentrated in a single commit (99.5%) while BugsInPy shows more scattered blame (58.4% single-commit).
Using the same LLM for all experiments including replicated baselines, HAFixAgent outperforms RepairAgent (+56.6%) and BIRCH-feedback (+47.1%) on Defects4J. Historical context further improves repair by +4.4% on Defects4J and +38.6% on BugsInPy, with the largest gains on SFMH bugs (+190.7% on BugsInPy). Under noisy fault localization, history provides increasing resilience, particularly on SFMH where non-history collapses to 0% while history maintains up to 56%. History does not significantly inflate agent cost on either benchmark, and the three complementary heuristics each solve distinct bug subsets.
Two promising directions for future work emerge from our findings. First, HAFixAgent currently uses an LLM judge to select a single blame commit; for BugsInPy bugs where 41.6% have multiple blame commits, combining context from several commits could further improve repair effectiveness. Second, evolving HAFixAgent into a multi-agent architecture (Rafi et al., 2024; Lin et al., 2025; Lee et al., 2024) with specialized History, Fixer, and Reviewer agents could improve repair of complex MFMH bugs where a single agent’s context window is easily overwhelmed.